User Tools

Site Tools


lbaops:lbamar2008:vt11g1holog

The log from experiment vt11g1 has been copied to:

ftp.atnf.csiro.au:/pub/people/vlbi/incoming/

ANTAB and UVFLG files have been copied to the same place.

Additional experiment notes:

eVLBI

observers: Brett Reid, Jamie McCallum

Comments follow:

2008.087.02:42:09.77:“ vt11gs 2008 HOBART H Ho

2008.087.02:52:31.54;”Coherence tone 841.4 MHz works

2008.087.02:53:17.51;“Agilent set at 14 GHz, for first LO of 8.8 GHz

2008.087.02:53:26.67;”2nd LO set to 777 MHz.

2008.087.02:53:54.81;“System temperatures ~ 500 Jy as expected.

2008.087.02:56:36.22;”DAS profile is VSOP.PRO

2008.087.04:53:06.61:“ vt11g1 2008 HOBART H Ho

2008.087.05:03:48.50;”Drives turned themselves off during a slew. Now on and slewing to source.

2008.087.09:46:27.05;“System coherent

Regards,

Brett

lbaops/lbamar2008/vt11g1holog.txt · Last modified: 2015/12/18 16:38 (external edit)